# SCOPE: Stochastic Cartographic Occupancy Prediction Engine for Uncertainty-Aware Dynamic Navigation

### v1

@@ IV-C Uncertainty Quantification
where the first term on RHS is the truncated normal distribution (with mean $\mu_{\rm tn}$, standard deviation $\sigma_{\rm tn}$, lower bound $a$, and upper bound $b$), the second RHS term is the skew-Cauchy distribution (with location parameter $\mu_{\rm sc}$, scale parameter $\sigma_{\rm sc}$, and skewness $\lambda$), and $w\in[0,1]$ is the mixture weight.
Let $\mathbf{\xi}=[w,a,b,\mu_{\rm tn},\sigma_{\rm tn},\lambda,\mu_{\rm sc},\sigma_{\rm sc}]$ be the vector of mixture model parameters. We use a least square optimization algorithm [67] to fit $\mathbf{\xi}$ to our predicted OGM cell bin data. Figure 3 shows the results of this fit. We see that the model achieves strong performance across a range of bins and timesteps, indicating that our methods works well.
The end result of our analysis above is a set 150 $\mathbf{\xi}$ vectors, one for each of the 15 $\hat{c}$ bins and each of the 10 time steps $T$. For each set of mixture model parameters, we can generate different types of distribution statistics (e.g., mean, median, mode, entropy, and random variate) and store these values in a lookup table. We call this lookup table the prediction uncertainty statistics lookup table.
Now that we have accurate models for the statistics of how each cell in the map will change over time, we can integrate that information into the SO-SCOPE model. To do this, we use the output of the distillation layer to get $\hat{c}_{t+T}$ for each cell in the OGM $\mathbf{o}_{t+T}$, as shown in the SO-SCOPE part of Fig. 2. We can then use the prediction uncertainty statistics lookup table to get statistics of that cell, or we can use the cumulative distribution function of (6) to draw samples $\tilde{c}_{t+T}$. This allows our SO-SCOPE model to account for uncertainty in the predicted future output using knowledge distilled from the SCOPE and SCOPE++ models to achieve comparable accuracy with significantly smaller computational effort and memory consumption.
@@ V OGM Prediction Experiments and Results
(a) OGM-Turtlebot2 dataset
(b) OGM-Jackal dataset
(c) OGM-Spot dataset
Figure 4: Illustration of the OGM dataset collection from three different robots.
@@ V-A Datasets
We used three publicaly available OGM datasets to evaluate our proposed prediction algorithms and baselines, one self-collected from the Gazebo simulator (i.e., OGM-Turtlebot2) [69] and two extracted from the real-world dataset SCAND [68] (i.e., OGM-Jackal and OGM-Spot).
OGM-Turtlebot2 Dataset: To train our predictor networks and baselines, we collected the OGM-Turtlebot2 dataset using the 3D robot-pedestrian interaction Gazebo simulator from our previous work [9, 11]. Specifically, a simulated Turtlebot2 equipped with a 2D Hokuyo UTM-30LX lidar navigates around an indoor environment with 34 moving pedestrians using random start points and goal points, as shown in Fig. 4a. The Turtlebot2 uses the dynamic window approach (DWA) planner [8] and has a maximum speed of 0.8 m/s. The pedestrians are driven by the social force model [70] with an average speed of 1.5 m/s. We collected the robot states $\{\mathbf{x},\mathbf{u}\}$ and raw lidar measurements $\mathbf{y}$ at a sampling rate of 10 Hz. We collected a total of 94,891 $\mathbf{d}=\{\mathbf{x},\mathbf{u},\mathbf{y}\}$ data tuples, dividing this into three separate subsets for training (67,000 tuples), validation during training (10,891 tuples), and final testing (17,000 tuples).
OGM-Jackal and OGM-Spot Datasets: Figure 4b shows the real-world outdoor environment at UT Austin and the Jackal robot used to collect the raw SCAND dataset to construct the OGM-Jackal dataset. Figure 4c shows the real-world indoor environment at UT Austin and the Spot robot used to collect the raw SCAND dataset to construct the OGM-Spot dataset. Note that the SCAND dataset was collected by humans manually operating the Jackal robot and the Spot robot around the indoor/outdoor environments at UT Austin. More details can be found in [68].
@@ V-B OGM Prediction
We compare our proposed SCOPE, SCOPE++, and SO-SCOPE algorithms with four deep learning-based baselines: ConvLSTM [5], DeepTracking [6], PhyDNet [7], and five ablation baselines: SCOPE-NEMC (SCOPE with No Ego-Motion Compensation module), SO-SCOPE-MEAN (use the mean value from the prediction uncertainty statistics lookup table as the OGM prediction), SO-SCOPE-MEDIAN (use the median from the lookup table), SO-SCOPE-MODE (use the mode from the lookup table), and SO-SCOPE-SAMPLE (draw a random sample from distribution). Note that all these networks were implemented using PyTorch framework [71] and trained using only our self-collected OGM-Turtlebot2 dataset.
@@ V-B1 Evaluation Metrics
We use three error metrics: 1) weighted mean square error (WMSE) [72] to evaluate absolute error between the ground truth and predicted maps, 2) structural similarity index measure (SSIM) [73] to measure the similarity between the two maps, and 3) optimal subpattern assignment metric (OSPA) [74] to measure the difference in the number and locations of distinct objects within the maps. Note that other OGM prediction works [14, 16, 17, 18, 15, 20, 21, 22, 25] only use the computer vision metrics (e.g., MSE, F1 Score, SSIM, and confusion matrices) to evaluate the quality of predicted OGMs. We are the first to evaluate the predicted OGMs from the perspective of multi-target tracking (i.e., OSPA distance), which we believe give a more physically meaningful evaluation than image quality as it accounts for the number of and locations of objects in the scene.
Figure 5: Evaluation pipeline for calculating OSPA error on predicted OGMs.
Figure 5 outlines our process to calculate the OSPA. First, we binarize the predicted and ground truth OGMs with an occupancy threshold $p_{\rm free}=0.3$, which is set by referencing the occ_thresh default parameter of 0.25 from the gmapping ROS package. Second, we use the density-based spatial clustering of applications with noise (DBSCAN) [75] algorithm to cluster the obstacle points in the OGMs. This yields estimated and ground truth sets of object locations in the scene. We then apply the OSPA error metric [74] between these object sets (with cutoff distance 10 cells, or 1 m).
@@ V-B2 Qualitative Results
Figure 6 and the accompanying multimedia illustrate the future OGM predictions generated by our proposed predictors and the baselines. We observe three interesting phenomena. First, the image-based baselines, especially the PhyDNet, generate blurry future predictions after 5 time steps, with only blurred shapes of static objects (i.e., walls) and missing dynamic objects (i.e., pedestrians). We believe that this is because these three baselines are deterministic models that use less expressive network architectures, only treat time series OGMs as images/video, and cannot capture and utilize the kinematics and dynamics of the robot itself, dynamic objects, and static objects. Second, the SCOPE++ with a local environment map has a sharper and more accurate surrounding scene geometry (i.e., right walls) than the SCOPE without a local environment map. This difference indicates that the local environment map for static objects is beneficial and plays a key role in predicting surrounding scene geometry. Third, our proposed software-optimized SO-SCOPE predictor is able to achieve clear and sharp OGM predictions similar to its “teacher” SCOPE predictor, which demonstrates the effectiveness of applying knowledge distillation techniques to optimize our SCOPE++ predictor.
Figure 6: A prediction showcase of seven OGM predictors tested on the OGM-Turtlebot2 dataset over the prediction horizon. The black area is the free space and the white area is the occupied space.
@@ V-B3 Quantitative Results
While achieving sharper predictions is encouraging, we also care about the quantitative measures of our proposed approaches. We test the OGM predictors on the OGM-Turtlebot2 test dataset, OGM-Jackal dataset, and OGM-Spot dataset by evaluating absolute error, structure similarity, and tracking accuracy. Recall that we only use the OGM-Turtlebot2 during training, so the other two datasets are entirely new and were generated by robots with different speeds and kinematics (i.e., Spot ambulates).
@@
To evaluate the absolute error of predicted OGMs, we calculate the average WMSE of predicted OGMs for the next 10 prediction time steps, outlined in Fig. 7. It can be seen that our proposed SCOPE family of methods with ego-motion compensation achieve significantly better average WMSE, SSIM, and OSPA than the SCOPE-NEMC ablation baseline without ego-motion compensation in all test datasets, which illustrates the effectiveness of ego-motion compensation. Furthermore, the average WMSEs of our proposed SCOPE family of methods at different prediction time steps are not significantly different from one another, but are all lower than the three image-based baselines (i.e., ConvLSTM [5], DeepTracking [6], and PhyDNet [7]). This trend holds across all three test datasets collected by different robots. This shows that the proposed SCOPE family of methods utilizing kinematic and dynamic information are able to predict more accurate OGMs than the state-of-the-art image-based approaches at different prediction time steps.
(a) OGM-Turtlebot2
(b) OGM-Jackal
(c) OGM-Spot
Figure 7: Average WMSE (lower is better) of 10 different prediction time steps for all tested methods on different datasets. Note that the average WMSE shadows for SCOPE-NEMC, SCOPE and SCOPE++ approaches are plotted with 95% confidence interval over 32 samples. For all curves, lower is better.
@@
To evaluate the structure similarity of predicted OGMs, we calculate the average SSIM of predicted OGMs for the next 10 prediction time steps, as shown in Fig. 8. Unlike general WSME, which cares about the absolute error of a single OGM cell, SSIM cares about the entire OGM structure and its scene geometry. Figure 8 shows three interesting phenomena. First, in all three different test datasets, the average SSIMs of our proposed SCOPE, SCOPE++, and SO-SCOPE predictors at different prediction time steps are significantly higher than that of other three image-based baselines and SCOPE-NEMC, which further illustrates the effectiveness of ego-motion compensation. Second, the average SSIMs of our software-optimized SO-SCOPE predictor and its statistical ablation baseline SO-SCOPE-MODE (i.e., using peak values as cell prediction values) are not significantly different (similar to that of the SCOPE++ predictor), but much higher than the other three statistical ablation baselines (i.e., SO-SCOPE-MEAN, SO-SCOPE-MEDIAN, SO-SCOPE-SAMPLE). It shows that the direct output of SO-SCOPE “student” network is able to represent the prediction information and the mode from our prediction uncertainty statistics lookup table is also a good statistic to represent the prediction information (compared to other statistics like mean, median, and random variate), which further supports the choice of using knowledge distillation techniques and uncertainty quantification to software optimize our original SCOPE/SCOPE++ predictors. Third, it is interesting that the average SSIM of our SCOPE++ with a local environment map is significantly higher than that of SCOPE without a local environment map in long-term predictions over multiple time steps. This indicates that the local environment map that takes into account static objects helps to predict the long-term future of the environment. This is reasonable because the local environment map describes the basic scene geometry and shape.
(a) OGM-Turtlebot2
(b) OGM-Jackal
(c) OGM-Spot
Figure 8: Average SSIM (higher is better) of 10 different prediction time steps for all tested methods on different datasets. Note that the average SSIM shadows for SCOPE-NEMC, SCOPE, and SCOPE++ approaches are plotted with 95% confidence interval over 32 samples. For all curves, higher is better.
@@
To further evaluate the tracking accuracy of predicted OGMs, we calculate the average OSPA of predicted OGMs for the next 10 prediction time steps, as shown in Fig. 9. While the average WMSE and SSIM are the evaluation metrics from computer vision, the average OSPA is from the multi-target tracking and is more suitable for evaluating the physical quality of OGMs.
@@
Figure 9 reveals that the average OSPA errors of our proposed SCOPE, SCOPE++, and SO-SCOPE predictors are significantly lower than the average OSPA errors of the other three image-based baselines and SCOPE-NEMC, especially in longer prediction time steps. This exciting result further demonstrates the preferential performance of our proposed motion-based methods (i.e., SCOPE, SCOPE++, and SO-SCOPE) in tracking or predicting environmental states (i.e., localization and cardinality) over other image-based methods (i.e., ConvLSTM [5], DeepTracking [6], and PhyDNet [7]). In addition, the average OSPA errors of our software-optimized SO-SCOPE predictor and its statistical ablation baselines are not significantly different from one another (similar to the WMSE metric), but are lower than the original SCOPE and SCOPE++ preditor (especially in the real-world OGM-Jackal and OGM-Spot dataset), which is reasonable and consistent with the argument in many knowledge distillation works [41, 65] that the “student” network can generalize better than the “teacher” network. However, the average OSPA errors of our proposed SCOPE and SCOPE++ predictors are almost the same, where the local environment map for static objects does not help SCOPE++ to reduce the OSPA error. We believe that this is because the local environment map only provide useful information for static objects while the OSPA metric is biased towards dynamic objects (e.g., pedestrians) since there tend to be more of them than than static objects (e.g., walls).
(a) OGM-Turtlebot2
(b) OGM-Jackal
(c) OGM-Spot
Figure 9: Average OSPA (lower is better) of 10 different prediction time steps for all tested methods on different datasets. Note that the average OSPA shadows for SCOPE-NEMC, SCOPE, and SCOPE++ approaches are plotted with 95% confidence interval over 32 samples. For all curves, lower is better.
Summary
Taken together, these meaningful results on the absolute error of a single occupancy grid cell, the structure similarity of overall scene geometry, and the tracking accuracy of multi-target localization and cardinality demonstrate that our proposed SCOPE, SCOPE++, and SO-SCOPE predictors with uncertainty estimates can predict sharper and more accurate long-term future states of the environment than image-based baselines.
V-C Computational Resource Utilization

### v2

@@ IV-C Uncertainty Quantification
where the first term on RHS is the truncated normal distribution (with mean $\mu_{\rm tn}$, standard deviation $\sigma_{\rm tn}$, lower bound $a$, and upper bound $b$) and the second RHS term is the skew-Cauchy distribution (with location parameter $\mu_{\rm sc}$, scale parameter $\sigma_{\rm sc}$, and skewness $\lambda$), and $w\in[0,1]$ is the mixture weight.
Let $\mathbf{\xi}=[w,a,b,\mu_{\rm tn},\sigma_{\rm tn},\lambda,\mu_{\rm sc},\sigma_{\rm sc}]$ be the vector of mixture model parameters. We use a least square optimization algorithm [69] to fit $\mathbf{\xi}$ to our predicted OGM cell bin data. Figure 3 shows the results of this fit. We see that the model achieves strong performance across a range of bins and timesteps, indicating that our methods work well.
The end result of our analysis above is a set of 150 $\mathbf{\xi}$ vectors, one for each of the 15 $\hat{c}$ bins and each of the 10 time steps $T$. For each set of mixture model parameters, we can generate different types of distribution statistics (e.g., mean, median, mode, entropy, and random variate) and store these values in a lookup table. We call this lookup table the prediction uncertainty statistics lookup table.
Now that we have accurate models for the statistics of how each cell in the map will change over time, we can integrate that information into the SO-SCOPE model. To do this, we use the output of the distillation layer to get $\hat{c}_{t+T}$ for each cell in the OGM $\mathbf{o}_{t+T}$, as shown in the SO-SCOPE part of Fig. 2. We can then use the uncertainty statistics lookup table to get statistics of that cell, or we can use the cumulative distribution function of (6) to draw samples $\tilde{c}_{t+T}$. This allows our SO-SCOPE model to account for uncertainty in the predicted future using knowledge distilled from the SCOPE and SCOPE++ models to achieve comparable accuracy with significantly smaller computational effort and memory consumption.
@@ V-A Datasets
We used three publicly available OGM datasets to evaluate our proposed prediction algorithms and baselines, one self-collected from the Gazebo simulator (i.e., OGM-Turtlebot2) [14] and two extracted from the real-world dataset SCAND [70] (i.e., OGM-Jackal and OGM-Spot).
V-A1 OGM-Turtlebot2 Dataset
To train our predictors and baselines, we collected the OGM-Turtlebot2 dataset using the 3D robot-pedestrian interaction Gazebo simulator from our previous work [11, 13]. We collected the robot states $\{\mathbf{x},\mathbf{u}\}$ and raw lidar measurements $\mathbf{y}$ at a sampling rate of 10 Hz. We collected a total of 94,891 $\mathbf{d}=\{\mathbf{x},\mathbf{u},\mathbf{y}\}$ data tuples, dividing this into three separate subsets for training (67,000 tuples), validation during training (10,891 tuples), and final testing (17,000 tuples). More dataset details can be found in [4].
V-A2 OGM-Jackal and OGM-Spot Datasets
The OGM-Jackal and OGM-Spot datasets were processed and created from the original SCAND dataset [70], which was collected by humans manually controlling the Jackal robot and the Spot robot around the indoor/outdoor flat environments at UT Austin. [^5] Force control for complex dynamic motions like the Spot robot is handled by the low-level controller, while the human motion control commands collected in the dataset are high-level 2D velocity control commands. More dataset processing details can be found in [4].
@@ V-B OGM Prediction
We compare our proposed SCOPE, SCOPE++, and SO-SCOPE algorithms with six deep learning-based baselines: ConvLSTM [5], DeepTracking [6], PhyDNet [7], SAAConvLSTM [8], TAAConvLSTM [8], and LOPR [9], and six ablation baselines: SCOPE-NEMC (SCOPE with No Ego-Motion Compensation module), SO-SCOPE-GTEMC (SO-SCOPE with Ground Truth of Ego-Motion Compensation module), SO-SCOPE-MEAN (use the mean value from the prediction uncertainty statistics lookup table as the OGM prediction), SO-SCOPE-MEDIAN (use the median from the lookup table), SO-SCOPE-MODE (use the mode from the lookup table), and SO-SCOPE-SAMPLE (draw a random sample from distribution)  [^6] Note that we ignore two ablation baselines of training the same SO-SCOPE network using dataset labels and using the SCOPE++ model as the “teacher” network, as they are both used to support the knowledge distillation theory and validated by previous researchers [67, 46, 68].. Note that SCOPE-NEMC is used to demonstrate the advantages of our proposed ego-motion compensation module and the network structure compared to other image-based baselines without ego-motion compensation. SO-SCOPE-GTEMC is used to demonstrate the upper-performance limit of the SO-SCOPE predictor with accurate ego-motion compensation. All these networks were implemented using PyTorch framework [71] and trained using only our self-collected OGM-Turtlebot2 dataset.
@@ V-B1 Evaluation Metrics
We use three metrics: 1) weighted mean square error (WMSE) [72] to evaluate absolute error between the ground truth and predicted maps, 2) structural similarity index measure (SSIM) [73] to measure the similarity between the two maps, and 3) optimal subpattern assignment metric (OSPA) [74] to measure the difference in the number and locations of distinct objects within the maps. Note that OGM prediction works [17, 19, 20, 8, 9, 18, 22, 23, 24, 27] only use the computer vision metrics (e.g., MSE, F1 Score, and SSIM) to evaluate the quality of predicted OGMs. Since a blurred prediction with uncertainty may perform better in some cases, we are the first to evaluate the predicted OGMs from the perspective of multi-target tracking (i.e., OSPA distance), which we believe gives a more physically meaningful evaluation than image quality as it accounts for the number of and locations of objects in the scene. More details can be found in [4].
@@ V-B2 Qualitative Results
Figure 4 and the accompanying multimedia illustrate the future OGM predictions generated by our proposed predictors and the baselines. We observe three interesting phenomena. First, the image-based baselines, especially the PhyDNet, generate blurry future predictions after 5-time steps, with only blurred shapes of static objects (i.e., walls) and missing dynamic objects (i.e., pedestrians). We believe this is because these six baselines are deterministic models that use less expressive network architectures, only treat time series OGMs as images/video, and cannot capture and utilize the kinematics and dynamics of the robot itself, dynamic objects, and static objects. Second, the SCOPE++ with a local environment map has a sharper and more accurate surrounding scene geometry (i.e., right walls) than the SCOPE without it. This difference indicates that the local environment map for static objects is beneficial and plays a key role in predicting surrounding scene geometry. Third, our proposed software-optimized SO-SCOPE can achieve clear and sharp OGM predictions similar to its “teacher” SCOPE, which demonstrates the effectiveness of applying knowledge distillation techniques to optimize our SCOPE/SCOPE++.
Figure 4: Prediction showcase of ten OGM predictors tested on the OGM-Turtlebot2 dataset. The black and white areas are free and occupied space respectively.
@@ V-B3 Quantitative Results
While achieving sharper predictions is encouraging, we also need quantitative measures. We test 14 OGM predictors on the OGM-Turtlebot2 test dataset, OGM-Jackal dataset, and OGM-Spot dataset and provide a comprehensive benchmark by evaluating absolute error, structure similarity, and tracking accuracy. Recall that we only use the OGM-Turtlebot2 during training, so the other two datasets are entirely new and were generated by robots with different speeds and kinematics (i.e., Spot ambulates).
@@
To evaluate the absolute error of predicted OGMs, we calculate the average WMSE of predicted OGMs for the next 10 prediction time steps, outlined in Figs. 5a, 5b and 5c. It can be seen that our proposed SCOPE family of methods with ego-motion compensation achieves significantly better average WMSE than the SCOPE-NEMC ablation baseline without ego-motion compensation in all test datasets, which illustrates the effectiveness of ego-motion compensation. The average WMSEs of our proposed SCOPE family of methods at different prediction time steps are not significantly different from one another, but are all lower than the six image-based baselines (i.e., ConvLSTM [5], DeepTracking [6], PhyDNet [7], SAAConvLSTM [8], TAAConvLSTM [8], and LOPR [9]). This trend holds across all three test datasets collected by different robots. This shows that the proposed SCOPE family of methods utilizing kinematic and dynamic information can predict more accurate OGMs than image-based approaches at different prediction time steps. Furthermore, the average WMSEs of the SO-SCOPE-GTEMC are much higher than that of the SO-SCOPE, which indicates that the best prediction performance can be achieved if our SO-SCOPE is equipped with an accurate customized ego-motion compensation algorithm for the robot instead of using only a general constant velocity motion model.
(a) WMSE: OGM-Turtlebot2
(b) WMSE: OGM-Jackal
(c) WMSE: OGM-Spot
(d) SSIM: OGM-Turtlebot2
(e) SSIM: OGM-Jackal
(f) SSIM: OGM-Spot
(g) OSPA: OGM-Turtlebot2
(h) OSPA: OGM-Jackal
(i) OSPA: OGM-Spot
Figure 5: WMSE (row 1), SSIM (row 2), and OSPA (row 3) for all tested methods on our 3 different datasets (columns). Each figure shows how the average value (over samples in the test datasets) of a given metric changes as a function of the prediction horizon, where lower is better for WMSE and OSPA and higher is better for SSIM. Lines for SCOPE-NEMC, SCOPE, and SCOPE++ include a 95% confidence interval over 32 samples drawn from the VAE module.
@@
To evaluate the structure similarity of predicted OGMs, we calculate the average SSIM of predicted OGMs for the next 10 prediction time steps, as shown in Figs. 5d, 5e and 5f. Unlike general WSME, which cares about the absolute error of a single OGM cell, SSIM cares about the entire OGM structure and its scene geometry. First, we can see that the average SSIMs of our proposed SCOPE, SCOPE++, and SO-SCOPE predictors at different prediction time steps are significantly higher than that of SCOPE-NEMC, and the average SSIMs of SCOPE-NEMC are almost higher than the other six image-based baselines, which further illustrates the effectiveness of ego-motion compensation and also the advantage of our network design. Second, the average SSIMs of our software-optimized SO-SCOPE predictor and its statistical ablation baseline SO-SCOPE-MODE (i.e., using peak values as cell prediction values) are not significantly different (similar to that of the SCOPE++ predictor), but much higher than the other three statistical ablation baselines (i.e., SO-SCOPE-MEAN, SO-SCOPE-MEDIAN, SO-SCOPE-SAMPLE). It shows that the direct output of SO-SCOPE “student” network can represent the prediction information and the mode from our prediction uncertainty statistics lookup table is also a good statistic to represent the prediction information (compared to other statistics like mean, median, and random variate), which further supports the choice of using knowledge distillation techniques and uncertainty quantification to software optimize our original SCOPE/SCOPE++ predictors. Third, it is interesting that the average SSIM of our SCOPE++ with a local environment map is significantly higher than that of SCOPE without a local environment map in long-term predictions over multiple time steps. This indicates that the local environment map that takes into account static objects helps to predict the long-term future of the environment. This is reasonable because the local environment map describes the basic scene geometry and shape. Finally, the highest average SSIMs of SO-SCOPE-GTEMC also show the upper limit of the performance of our SO-SCOPE.
(a) Inaccurate robot motion compensation
(b) Occlusion or sudden appearance
Figure 6: Two typical prediction failure cases of the SCOPE series predictors (SO-SCOPE as the example). The red bounding boxes highlight prediction errors.
TABLE II: Inference Speed, Model Size, and Memory Usage of different OGM predictors
Models ConvLSTM [5] PhyDNet [7] DeepTracking [6] SAAConvLSTM [8] TAAConvLSTM [8] LOPR [9] SCOPE SCOPE++ SO-SCOPE
Inference Speed (FPS) 2.95 4.66 5.32 0.39 0.39 1.16 23.29 10.68 34.75
Model Size (MB) 12.44 37.17 0.95 21.64 22.02 1610.48 8.84 8.85 1.80
Memory Usage (GB) 0.70 0.71 0.63 1.23 1.23 5.00 0.66 0.66 0.66
@@
To further evaluate the tracking accuracy of predicted OGMs, we calculate the average OSPA of predicted OGMs for the next 10 prediction time steps, as shown in Figs. 5g, 5h and 5i. While the average WMSE and SSIM are the evaluation metrics from computer vision, the average OSPA is from the multi-target tracking and is more suitable for evaluating the physical quality of OGMs. Figures 5g, 5h and 5i reveals that the average OSPA errors of our proposed SCOPE family of predictors are significantly lower than that of SCOPE-NEMC, and the average OSPA errors of SCOPE-NEMC are almost lower than that of other image-based baselines. This exciting result further demonstrates the preferential performance of our proposed motion-based methods (i.e., SCOPE, SCOPE++, and SO-SCOPE) in tracking or predicting environmental states (i.e., localization and cardinality) outperform SCOPE-NEMC, and while SCOPE-NEMC almost outperforms other image-based baselines (similar to the SSIM metric). Furthermore, the average OSPA errors of our software-optimized SO-SCOPE predictor and its statistical ablation baselines are not significantly different from one another (similar to the WMSE metric), but are lower than the original SCOPE and SCOPE++ predictors (especially in the real-world OGM-Jackal and OGM-Spot datasets), which is attributed to the use of knowledge distillation techniques to distill stochastic neural network models (i.e., data augmentation). It is reasonable and consistent with the argument in many knowledge distillation works [43, 67] that the “student” network can generalize better than the “teacher” network. However, the average OSPA errors of our proposed SCOPE and SCOPE++ predictors are almost the same, where the local environment map for static objects does not help SCOPE++ to reduce the OSPA error. We believe that this is because the local environment map only provides useful information for static objects while the OSPA metric is biased towards dynamic objects (e.g., pedestrians) since there tend to be more of them than static objects (e.g., walls). Finally, the highest average OSPAs of SO-SCOPE-GTEMC also show the upper limit of the performance of our SO-SCOPE.
V-C Failure Case Analysis
Although our SCOPE series demonstrates its accurate and robust future state prediction ability, there are still two typical failure cases. The first failure case is when the robot moves erratically or rotates rapidly and the robot motion compensation is inaccurate, as shown in Fig. 6a. We can easily see that when the robot suddenly rotates at the 6th prediction time step, the SO-SCOPE predictor of our default constant velocity motion model fails to predict the later future states. This is because our general constant velocity motion model fails to predict the robot’s future motion, which is the key to OGM prediction. To support our analysis, we also provide prediction results for the SO-SCOPE-GTEMC (which shows an upper bound on the performance of SO-SCOPE with ground truth motions). From the last row of Fig. 6a, we can see that if we can accurately compensate for the robot’s motion, our SO-SCOPE can still provide accurate future predictions after the 6th prediction timestep. Therefore, this failure issue can be easily alleviated by using more accurate robot motion models specific to the robots the researchers are using, as described in Sec. III-C1.
The second failure case is when the robot’s field of view (FOV) is blocked by a large obstacle or objects such as pedestrians suddenly appear in the robot’s FOV, as shown in Fig. 6b. We can see that when two pedestrians suddenly appear in the robot’s FOV at the 6th prediction time step, our SO-SCOPE still misses the future states of these two pedestrians even with accurate motion compensation (i.e., SO-SCOPE-GTEMC). This is because there is not enough information about the object’s motion to feed into the neural network predictor, which is also important for OGM prediction. We believe that adding additional randomly sampled particles to the input OGMs or providing a longer OGM history can mitigate this occlusion or sudden appearance failure case. Exploring the OGM prediction problem in occluded or sudden appearance scenes will be our future work.
V-D Computational Resource Utilization